### v1

@@ 3.3 Vertical structure of the upwelling cell
Figure 6: Cross-shore transects of $VT_{\mbox{up-seas}}$ (color-filled black contours) for the Canary (top) and Benguela (bottom) upwelling system, within each sub-region (column). $VT_{\mbox{up-seas}}$ corresponds to an along-shore seasonal average over the entire historical period ($1995-2014$). The three-month seasonal avarage is adapted to each sub-region to only consider months most favourable to upwelling. Contours of potential temperature (light grey) and mixed layer depth (solid white) are superimposed. Line plots above each transect show cross-shore profiles of Sea Surface Temperature (solid black), offshore Ekman transport (solid red)and $VT_{m}$ (blue). The model shown here is the HadGEM3-GC3.1-MM model.
@@
After validating the seasonal cycle of the indices, we choose the most intense upwelling season for each of the subregions to consider the vertical structure of the upwelling cell in each subregion. The resulting upwelling cells derived from the HadGEM3-GC3.1-MM model can be seen in Figure 6. We display above each cell the Ekman transport, the $VT_{m}$ and the SST. Looking at the upwelling cell, some features are similar to the annual means over the whole region shown in Figure 2. However, the different subregions also show differences in their vertical structures. In the Canary region, maxima are found closer to the coast in the NCUC and the CCUC, while the SCUC shows a higher zonal extent. This corresponds to maximal wind stress intensities close to the coast in the northern and central part, while in the south maxima are found further offshore. The highest upwelling rates are found in the SCUC, which has been described to be the most productive region in the Canary upwelling system (Sylla et al., 2019). In the Benguela upwelling the highest upwelling rates and vertical extent can be found in the CBUC, which is where the Lüderitz cell, the most intense upwelling cell in the Benguela system is located. This is in agreement with Chang et al. (2023). The SBUC shows only slightly lower upwelling rates and vertical extent, even though the transport calculated from wind is much weaker. The maximum in the SBUC is found further from the coast. This is mirrored in the line plots above the cell plots, where maximum Ekman transport calculated from meridional wind stress and maximum $VT_{m}$ peak at the same distance from the coast. As the Ekman transport resulting from wind stress forcing is the main driver of upwelling, this is an expected outcome, reflecting the ocean response to the atmospheric forcing. The difference between the Ekman transport at the coast and the Ekman transport offshore reflects the importance of the zonal gradient of the meridional wind stress. We have verified that this term is the most important component of the wind stress curl.
@@
The differences between the subregions in Figure 6 can be found in the vertical extent, the distance to the coast of the maximum and the depth of the maximum. They are influenced by static features, such as the coastal shape and shelf structure (Benazzouz et al., 2014). This might for example be a factor in the maxima occuring further offshore in the southern regions of both upwellings, as they are characterized by a wider, shallow continental shelf. Dynamical features also play a role, such as the strength and extent of the meridional wind stress forcing, and the Wind Stress Curl (WSC), which have an impact on the amount and depth from which water is upwelled (Capet et al., 2004). For future work, a closer analysis of the coastal bathymetry, coastline shape and WSC generated Ekman pumping on a regional scale in comparison with the divergence of Ekman transport could help explain differences between the depth from which water is upwelled. Regarding the relation between the Ekman Index and the Vertical Transport, the closer the cell is located towards the equator, the larger the difference becomes between the two indices, as already noted in section 3.2. This can have an important implication on upwelling processes as especially geostrophic transport has been found to have a counteracting or enhancing effect on upwelling through convergence or divergence (Veitch et al., 2010) and could play a high role in controlling the future of upwelling regions (Sylla et al., 2019; Jing et al., 2023). Jing et al. (2023) found especially high differences between their wind-based index and vertical transport based index in the Atlantic Upwelling Systems, which they ascribe to downwelling caused by geostrophic flows. They further state that analysing changes in the geostrophic flow could therefore help in determining future changes. As in Fig. 6 the differences between the $EI$ and the $VT_{m}$ are highest in the equatorward portions, our results suggest that regional differences of the impact of the ocean circulation within the two Atlantic upwellings could also be an important factor. In the Benguela upwelling, Veitch et al. (2010) have found an inhibition of upwelling in the northern part due to geostrophic convergence and an enhancement of upwelling in the southern part due to geostrophic divergence. The geostrophic convergence in the tropical parts of the upwelling regions is ascribed to the eastward equatorial currents of the Angola Dome in the Benguela and the Guinea Dome in the Canary upwelling system (Marchesiello and Estrade, 2010). This could hence explain the patterns observed in Fig. 6.
@@
We discuss the main differences between the models briefly (the figures for the remaining four models can be found in the supplementary material, Figures S.1-S.4). The vertical structure of the cells is comparable across models and resolutions. Consistent with higher wind stress as described in section 3.2, in general vertical transport is higher in the HadGEM models (Fig. 6, Fig. S.1 and S.2) than the CNRM ones (Fig. S.3 and S.4) as has been already observed in the seasonal cycle in Fig.4 and 5. In the Canary upwelling the influence of resolution is contrasted between the two models. For the HadGEM model vertical transport is weaker in all three subregions with higher resolution, while in the CNRM models transport is also weaker in the NCUC and SCUC in the high resolution model, but stronger in the CCUC. In the Benguela, in the HadGEM models the highest intensities in the CBUC and SBUC occur at the ¼ ° resolution. For CNRM, the transport is higher in the low resolution model. This behaviour is different from the CESM model used in Chang et al. (2023), where the lower resolution model shows lower transport in all four major upwelling regions. This indicates that both model and resolution can play an important role in the representation of fine scale differences in upwelling vertical water transport. This has also been investigated in a recent study by Sylla et al. (2022) analysing CMIP6 model performance in the Canary upwelling system, which has shown that an efficient method to have an improvement in upwelling representation is increasing both oceanic and atmospheric resolution. However, they also highlight the importance of model parametrization, tuning, spin-up period length and representation of processes.
@@ 3.4 Linear trends over the 21st century
Figure 7: 1995-2100 time series of yearly averaged $VTI$ in the CCUC subregion computed the HadGEM3-GC3.1-HH (solid red) and the HadGEM3-GC3.1-MM (solid black) simulations. The trend lines (dashed) are calculated using linear regression. While HadGEM3-GC3.1-HH shows a trend that is not significant with a p-value of 0.916, the HadGEM3-GC3.1-MM model shows a trend significantly different from 0 with a p-value of 0.006.
@@
Trends per century are calculated within each sub-region for the upwelling index $VTI$ presented in section 2.3. The trends are calculated over the 1995-2100 period from a linear regression performed on the yearly average values of the upwelling indices. The 90% confidence interval is computed from the student distribution with a number of degrees of freedom chosen after dividing the length of the time series by the corresponding lag associated with the 0.5 crossing of the auto-correlation function ($\simeq 12$). Note that the HadGEM3-GC3.1-HH simulation is excluded from our analysis because it only covers the 1995-2050 period. Some tests have been done on the definition of the depth and time scale of the index for the trend calculation. They have shown that for achieving statistically significant results, annual means provide better results than seasonal means. A constant depth is chosen to provide comparability with previous studies. The depth of 50 m having provided results closest to the Ekman index and showing the most significance, this is used for the calculation of the index for the future.
@@
The analysis of the computed trends allows us to revisit the two common hypothesis regarding the future of upwelling under global warming. Under the Bakun hypothesis (Bakun, 1990), an overall intensification of wind would occur and result in increased upwelling. Under the hypothesis of poleward migration however, we would expect upwelling intensities to increase (decrease) in the poleward (equatorward) subregions (Rykaczewski et al., 2015; Bograd et al., 2023). Using subregions allows us to assess the regional difference in trends, which is an important factor regarding the poleward migration hypothesis. In the Canary system a significant increase occurs in upwelling transport in the northern and central cell in both 1/4 ° resolution models and in the CCUC also for the CNRM-CM6-1 model. In the SCUC the models show less significant results and conflicting signs between the models. In the Benguela region both HadGEM models show a significant decrease in the northern cell, whereas in the central and southern region trends differ in sign between models and resolutions and do not pass the significance test. Since in the Benguela region we find some significant decrease in the NBUC, we can reject the Bakun hypothesis. And since not all our results are significant for all the subregions, we can neither accept nor reject the poleward migration hypothesis. Our results are however in support of the finding that trends differ regionally. This also highlights the impact of the exact region that is chosen for studies. In two recent works presented by Chang et al. (2023) and Jing et al. (2023) using the high resolution model Community Earth System Model (CESM) from the CMIP6 project and similar methods for assessing vertical transport, trends of conflicting signs have been found.
Figure 8: Trend per century in the Vertical Transport Index calculated over period 1995-2100 from historical simulation and SSP5 – 8.5 runs for each upwelling cell in the (a) Canary and (b) Benguela Upwelling region over annual means for the medium and low resolution models. The error bar shows the 90% confidence interval.
@@
In the Canary region our results are congruent with the results of Chang et al. (2023) and Jing et al. (2023). Chang et al. (2023) use the region from 36°N to 15°N, while Jing et al. (2023) look at the region between 32°N and 22°N. This covers mostly the NCUC and CCUC in our work. Overall positive trends are found in both studies, especially in the central part. This is also the case for our results. The SCUC shows contradicting trends in our work, this region is only partly considered (the northern section) in Chang et al. (2023) and excluded from the region chosen in Jing et al. (2023). In the Benguela region our results suggest a decrease in the NBUC and contradictory trends between the models in the CBUC and SBUC. Jing et al. (2023) find overall negative trends in upwelling transport in the Benguela region, while Chang et al. (2023) find an increase in upwelling transport. The latitudinal extent of the region is definitely an important factor. Jing et al. (2023) define their region as the area from 30°S to 10°S and Chang et al. (2023) from 32°S to 22°S. However, looking at the latitudinal distribution of the trends shows that even where their study regions overlap in the central Benguela, different results are obtained: positive trend in Chang et al. (2023) and negative trend in Jing et al. (2023). This can be explained by the two studies using a different longitudinal extent for their calculation, 50 km offshore for Chang et al. (2023) and 200 km offshore for Jing et al. (2023). The results differ because the vertical transport trends are opposite close to the coast compared with 100 km offshore (Ping Chang, personal communication). This highlights that mapping the vertical transport as a function of depth and distance to the coast, as done in this work, is a key factor when analysing possible future trends as they might differ locally. Regarding the relation between the wind-derived and the vertical transport derived index, Jing et al. (2023) find strong significant trends in their vertical transport index for both the CESM-H ensemble mean and the CMIP6 ensemble mean. For the Ekman index they find lower significant trends in the CESM-H ensemble mean, and very low, non-significant results in the CMIP6 ensemble mean. Our results in Figure 8 and S.10 show similar behaviour in significance.
@@ 5 Acknowledgments
6
@@
Figure S.1: Cross-shore transects of $VT_{\mbox{up-seas}}$ (color-filled black contours) for the Canary (top) and Benguela (bottom) upwelling system, within each sub-region (column). $VT_{\mbox{up-seas}}$ corresponds to an along-shore seasonal average over the entire historical period ($1995-2014$). The three-month seasonal avarage is adapted to each sub-region to only consider months most favourable to upwelling. Contours of potential temperature (light grey) and mixed layer depth (solid white) are superimposed. Line plots above each transect show cross-shore profiles of Sea Surface Temperature (solid black), offshore Ekman transport (solid red)and $VT_{m}$ (blue). The model shown here is the HadGEM3-GC3.1-HH model.
@@
Figure S.2: Cross-shore transects of $VT_{\mbox{up-seas}}$ (color-filled black contours) for the Canary (top) and Benguela (bottom) upwelling system, within each sub-region (column). $VT_{\mbox{up-seas}}$ corresponds to an along-shore seasonal average over the entire historical period ($1995-2014$). The three-month seasonal avarage is adapted to each sub-region to only consider months most favourable to upwelling. Contours of potential temperature (light grey) and mixed layer depth (solid white) are superimposed. Line plots above each transect show cross-shore profiles of Sea Surface Temperature (solid black), offshore Ekman transport (solid red)and $VT_{m}$ (blue). The model shown here is the HadGEM3-GC3.1-LL model.
@@
Figure S.3: Cross-shore transects of $VT_{\mbox{up-seas}}$ (color-filled black contours) for the Canary (top) and Benguela (bottom) upwelling system, within each sub-region (column). $VT_{\mbox{up-seas}}$ corresponds to an along-shore seasonal average over the entire historical period ($1995-2014$). The three-month seasonal avarage is adapted to each sub-region to only consider months most favourable to upwelling. Contours of potential temperature (light grey) and mixed layer depth (solid white) are superimposed. Line plots above each transect show cross-shore profiles of Sea Surface Temperature (solid black), offshore Ekman transport (solid red)and $VT_{m}$ (blue). The model shown here is the CNRM-CM6-1-HR model.
@@
Figure S.4: Cross-shore transects of $VT_{\mbox{up-seas}}$ (color-filled black contours) for the Canary (top) and Benguela (bottom) upwelling system, within each sub-region (column). $VT_{\mbox{up-seas}}$ corresponds to an along-shore seasonal average over the entire historical period ($1995-2014$). The three-month seasonal avarage is adapted to each sub-region to only consider months most favourable to upwelling. Contours of potential temperature (light grey) and mixed layer depth (solid white) are superimposed. Line plots above each transect show cross-shore profiles of Sea Surface Temperature (solid black), offshore Ekman transport (solid red)and $VT_{m}$ (blue). The model shown here is the CNRM-CM6-1 model.
@@
Figure S.5: Cross-shore transects of $VT_{\mbox{up-seas}}$ (color-filled black contours) for the Canary (top) and Benguela (bottom) upwelling system, within each sub-region (column). $VT_{\mbox{up-seas}}$ corresponds to an along-shore seasonal average over the entire future period ($2031-2050$). The three-month seasonal avarage is adapted to each sub-region to only consider months most favourable to upwelling. Contours of potential temperature (light grey) and mixed layer depth (solid white) are superimposed. Line plots above each transect show cross-shore profiles of Sea Surface Temperature (solid black), offshore Ekman transport (solid red)and $VT_{m}$ (blue). The model shown here is the HadGEM3-GC3.1-HH model.
@@
Figure S.6: Cross-shore transects of $VT_{\mbox{up-seas}}$ (color-filled black contours) for the Canary (top) and Benguela (bottom) upwelling system, within each sub-region (column). $VT_{\mbox{up-seas}}$ corresponds to an along-shore seasonal average over the entire future period ($2081-2100$). The three-month seasonal avarage is adapted to each sub-region to only consider months most favourable to upwelling. Contours of potential temperature (light grey) and mixed layer depth (solid white) are superimposed. Line plots above each transect show cross-shore profiles of Sea Surface Temperature (solid black), offshore Ekman transport (solid red)and $VT_{m}$ (blue). The model shown here is the HadGEM3-GC3.1-MM model.
@@
Figure S.7: Cross-shore transects of $VT_{\mbox{up-seas}}$ (color-filled black contours) for the Canary (top) and Benguela (bottom) upwelling system, within each sub-region (column). $VT_{\mbox{up-seas}}$ corresponds to an along-shore seasonal average over the entire future period ($2081-2100$). The three-month seasonal avarage is adapted to each sub-region to only consider months most favourable to upwelling. Contours of potential temperature (light grey) and mixed layer depth (solid white) are superimposed. Line plots above each transect show cross-shore profiles of Sea Surface Temperature (solid black), offshore Ekman transport (solid red)and $VT_{m}$ (blue). The model shown here is the HadGEM3-GC3.1-LL model.
@@
Figure S.8: Cross-shore transects of $VT_{\mbox{up-seas}}$ (color-filled black contours) for the Canary (top) and Benguela (bottom) upwelling system, within each sub-region (column). $VT_{\mbox{up-seas}}$ corresponds to an along-shore seasonal average over the entire future period ($2081-2100$). The three-month seasonal avarage is adapted to each sub-region to only consider months most favourable to upwelling. Contours of potential temperature (light grey) and mixed layer depth (solid white) are superimposed. Line plots above each transect show cross-shore profiles of Sea Surface Temperature (solid black), offshore Ekman transport (solid red)and $VT_{m}$ (blue). The model shown here is the CNRM-CM6-1-HR model.

### v2

@@ 3.3 Vertical structure of the upwelling cell
After validating the seasonal cycle of the indices, we choose the most intense upwelling season for each of the subregions to consider the vertical structure of the upwelling cell in each subregion. The resulting upwelling cells derived from the HadGEM3-GC3.1-MM model can be seen in Fig. 6. We display above each cell the Ekman transport, the $VT_{m}$ and the SST. Looking at the upwelling cell, some features are similar to the annual means over the whole region shown in Fig. 2. However, the different subregions also show differences in their vertical structures. In the Canary region, maxima are found closer to the coast in the NCUC and the CCUC, while the SCUC shows a higher zonal extent. This corresponds to maximal wind stress intensities close to the coast in the northern and central parts, while in the south maxima are found farther offshore. The highest upwelling rates are found in the SCUC, which has been described to be the most productive region in the Canary Upwelling system (Sylla et al., 2019). In the Benguela upwelling the highest upwelling rates and vertical extent can be found in the CBUC, which is where the Lüderitz cell, the most intense upwelling cell in the Benguela system, is located. This is in agreement with Chang et al. (2023). The SBUC shows only slightly lower upwelling rates and vertical extent, even though the transport calculated from wind is much weaker. The maximum in the SBUC is found farther from the coast. This is mirrored in the line plots above the cell plots, where maximum Ekman transport calculated from meridional wind stress and maximum $VT_{m}$ peak at the same distance from the coast. As the Ekman transport resulting from wind stress forcing is the main driver of upwelling, this is an expected outcome, reflecting the ocean response to the atmospheric forcing. The difference between the Ekman transport at the coast and the Ekman transport offshore reflects the importance of the zonal gradient of the meridional wind stress. We have verified that this term is the most important component of the wind stress curl.
@@
The differences between the subregions in Fig. 6 can be found in the vertical extent, the distance to the coast of the maximum values and the depth of the maximum values. They are influenced by static features, such as the coastal shape and shelf structure (Benazzouz et al., 2014). This might, for example, be a factor in the maxima occuring farther offshore in the southern regions of both upwellings, as they are characterized by a wider, shallow continental shelf. Dynamical features also play a role, such as the strength and extent of the meridional wind stress forcing, and the Wind Stress Curl (WSC), which have an impact on the amount and depth from which water is upwelled (Capet et al., 2004). For future work, a closer analysis of the coastal bathymetry, coastline shape and WSC generated Ekman pumping on a regional scale in comparison with the divergence of Ekman transport could help explain differences between the depth from which water is upwelled. Regarding the relation between the Ekman Index and the Vertical Transport, the closer the cell is located towards the equator, the larger the difference becomes between the two indices, as already noted in section 3.2. This can have an important implication on upwelling processes as especially geostrophic transport has been found to have a counteracting or enhancing effect on upwelling through convergence or divergence (Veitch et al., 2010) and could play a large role in controlling the future of upwelling regions (Sylla et al., 2019; Jing et al., 2023). Jing et al. (2023) found especially large differences between their wind-based index and vertical transport based index in the Atlantic Upwelling Systems, which they ascribe to downwelling caused by geostrophic flows. They further state that analysing changes in the geostrophic flow could therefore help in determining future changes. As in Fig. 6, the differences between the $EI$ and the $VT_{m}$ are highest in the equatorward portions, suggesting that regional differences of the impact of the ocean circulation within the two Atlantic upwellings could also be an important factor. In the Benguela upwelling, Veitch et al. (2010) have found an inhibition of upwelling in the northern part due to geostrophic convergence and an enhancement of upwelling in the southern part due to geostrophic divergence. The geostrophic convergence in the tropical parts of the upwelling regions is ascribed to the eastward equatorial currents of the Angola Dome in the Benguela and the Guinea Dome in the Canary Upwelling system (Marchesiello and Estrade, 2010). This could explain the patterns observed in Fig. 6.
@@
We discuss the main differences between the models briefly (the figures for the remaining four models can be found in the supplementary material, Figs. S.1-S.4). The vertical structure of the cells is comparable across models and resolutions. Consistent with higher wind stress as described in section 3.2, in general vertical transport is higher in the HadGEM models (Figs. 6, S.1 and S.2) than the CNRM ones (Figs. S.3 and S.4) as has been already observed in the seasonal cycle in Figs. 4 and 5. In the Canary upwelling, the influence of resolution is contrasted between the two models. For the HadGEM model, vertical transport is weaker in all three subregions with higher resolution, while in the CNRM models, transport is also weaker in the NCUC and SCUC in the high resolution model, but stronger in the CCUC. In the Benguela, in the HadGEM models the highest intensities in the CBUC and SBUC occur at the ¼ ° resolution. For CNRM, the transport is higher in the low resolution model. This behaviour is different from the CESM model used in Chang et al. (2023), where the lower resolution model shows lower transport in all four major upwelling regions. This indicates that both model and resolution can play an important role in the representation of fine scale differences in upwelling vertical water transport. This has also been investigated in a recent study by Sylla et al. (2022) analysing CMIP6 model performance in the Canary upwelling system, which has shown that an efficient method to have an improvement in upwelling representation is increasing both oceanic and atmospheric resolution. However, they also highlight the importance of model parametrization, tuning, spin-up period length and representation of processes.
@@ 3.4 Linear trends over the 21st century
Trends per century are calculated within each sub-region for the upwelling index $VTI$ presented in section 2.3. The trends are calculated over the 1995-2100 period from a linear regression performed on the yearly average values of the upwelling indices. The 90% confidence interval is computed from the student distribution with a number of degrees of freedom chosen after dividing the length of the time series by the corresponding lag associated with the 0.5 crossing of the auto-correlation function ($\simeq 12$ years). Note that the HadGEM3-GC3.1-HH simulation is excluded from our analysis because it only covers the 1995-2050 period. Some tests have been done on the definition of the depth and time scale of the index for the trend calculation. They have shown that for achieving statistically significant results, annual means provide better results than seasonal means. A constant depth is chosen to provide comparability with previous studies. The depth of 50 m having provided results closest to the Ekman index and showing the most significance, this is used for the calculation of the index for the future.
@@
The analysis of the computed trends allows us to revisit the two common hypothesis regarding the future of upwelling under global warming. Under the Bakun hypothesis (Bakun, 1990), an overall intensification of wind would occur and result in increased upwelling. Under the hypothesis of poleward migration however, we would expect upwelling intensities to increase (decrease) in the poleward (equatorward) subregions (Rykaczewski et al., 2015; Bograd et al., 2023). Using subregions allows us to assess the regional difference in trends, which is an important factor regarding the poleward migration hypothesis. In the Canary system, a significant increase occurs in upwelling transport in the northern and central cell in both 1/4° resolution models and in the CCUC also for the CNRM-CM6-1 model. In the SCUC, the models show less significant results and conflicting signs between the models. In the Benguela region both HadGEM models show a significant decrease in the northern cell, whereas in the central and southern region, trends differ in sign between models and resolutions and do not pass the significance test. Since in the Benguela region we find some significant decrease in the NBUC, we can reject the Bakun hypothesis. Since not all our results are significant for all the subregions, we can neither accept nor reject the poleward migration hypothesis. Our results, however, support the finding that trends differ regionally. This highlights the sensitivity of the results to the choice of the exact region chosen for the analysis. In two recent works presented by Chang et al. (2023) and Jing et al. (2023) using the high resolution model Community Earth System Model (CESM) from the CMIP6 project and similar methods for assessing vertical transport, trends of conflicting signs have been found.
@@
In the Canary region, our results are congruent with the results of Chang et al. (2023) and Jing et al. (2023). Chang et al. (2023) use the region from 36°N to 15°N, while Jing et al. (2023) look at the region between 32°N and 22°N. This covers mostly the NCUC and CCUC in our work. Overall positive trends are found in both studies, especially in the central part. This is also the case for our results. The SCUC shows contradicting trends in our work, this region is only partly considered (the northern section) in Chang et al. (2023) and excluded from the region chosen in Jing et al. (2023). In the Benguela region, our results suggest a decrease in the NBUC and contradictory trends between the models in the CBUC and SBUC. Jing et al. (2023) find overall negative trends in upwelling transport in the Benguela region, while Chang et al. (2023) find an increase in upwelling transport. The latitudinal extent of the region is definitely an important factor. Jing et al. (2023) define their region as the area from 30°S to 10°S and Chang et al. (2023) from 32°S to 22°S. However, looking at the latitudinal distribution of the trends shows that even where their study regions overlap in the central Benguela, different results are obtained: positive trend in Chang et al. (2023) and negative trend in Jing et al. (2023). This can be explained by the two studies using a different longitudinal extent for their calculation, 50 km offshore for Chang et al. (2023) and 200 km offshore for Jing et al. (2023). The results differ because the vertical transport trends are opposite close to the coast compared with 100 km offshore (Ping Chang, personal communication). This highlights that mapping the vertical transport as a function of depth and distance to the coast, as done in this work, is a key factor when analysing possible future trends as they might differ locally. Regarding the relation between the wind-derived and the vertical transport derived index, Jing et al. (2023) find strong significant trends in their vertical transport index for both the CESM-H ensemble mean and the CMIP6 ensemble mean. For the Ekman index they find lower significant trends in the CESM-H ensemble mean, and very low, non-significant results in the CMIP6 ensemble mean. Our results in Figs. 8 and S.10 show similar behaviour in significance.
@@ 5 Acknowledgments
(a) Refer to caption
(b) Refer to caption
(c) Refer to caption
(d) Refer to caption
Fig. 1: Seasonal mean Sea Surface Temperature from CCI Satellite data averaged over the historical 1995-2014 period for the Canary Upwelling region in boreal winter (a) and summer (b), and the Benguela Upwelling region in boreal winter (c) and summer (d). Black boxes show the subregions used for calculating indices.
Fig. 2: Annual mean vertical transport $VT$ (color-filled black contours in m2 s-1) averaged in time over the historical 1995-2014 period and in latitude over the entire region retained for the: (a) Canary Upwelling; (b) Benguela Upwelling. Contours of potential temperature (solid grey in C) and mixed layer depth (solid white) are superimposed. The model shown here is HadGEM3-GC3.1-HH.
Fig. 3: Annual mean SST bias showed by the CMIP6 models over the 1995-2014 period in the Atlantic EBUS: (a) Canary; (b) Benguela. $SST_{\mbox{bias}}=SST_{\mbox{CMIP6}}-SST_{\mbox{CCI}}$.
Fig. 4: Mean monthly upwelling indices over the time period 1995-2014 and their standard deviation for the three subregions in the Canary Upwelling system for the five CMIP6 models, the SST observations from CCI (Panels a,d,g) and the ERA5 Reanalysis wind stress (Panels b,e,h).
Fig. 5: Mean monthly upwelling indices over the time period 1995-2014 and their standard deviation for the three subregions in the Benguela Upwelling system for the five CMIP6 models, the SST observations from CCI (Panels a,d,g) and the ERA5 Reanalysis wind stress (Panels b,e,h).
Fig. 6: Cross-shore transects of $VT_{\mbox{up-seas}}$ (color-filled black contours) for the Canary (top) and Benguela (bottom) upwelling system, within each sub-region (column). $VT_{\mbox{up-seas}}$ corresponds to an along-shore seasonal average over the entire historical period ($1995-2014$). The three-month seasonal average is adapted to each sub-region to only consider months most favourable to upwelling. Contours of potential temperature (light grey) and mixed layer depth (solid white) are superimposed. Line plots above each transect show cross-shore profiles of Sea Surface Temperature (solid black), offshore Ekman transport (solid red)and $VT_{m}$ (blue). The model shown here is the HadGEM3-GC3.1-MM model.
Fig. 7: 1995-2100 time series of yearly averaged $VTI$ in the CCUC subregion computed from the HadGEM3-GC3.1-HH (solid red) and the HadGEM3-GC3.1-MM (solid black) simulations. The trend lines (dashed) are calculated using linear regression. While HadGEM3-GC3.1-HH shows a trend that is not significant with a p-value of 0.916, the HadGEM3-GC3.1-MM model shows a trend significantly different from 0 with a p-value of 0.006.
Fig. 8: Trend per century in the Vertical Transport Index calculated over period 1995-2100 from historical simulation and SSP5 – 8.5 runs for each upwelling cell in the (a) Canary and (b) Benguela Upwelling System over annual means for the medium and low resolution models. The error bar shows the 90% confidence interval.
Table 1: CMIP6 models used in this work, their horizontal resolution and number of vertical layers used for the atmospheric and oceanic component (Gutiérrez et al., 2021).
HadGEM3-GC31-LL 140km, 85L 70km, 75L ([dataset] Ridley et al., 2018)
HadGEM3-GC31-MM 60km, 85L 20km, 75L ([dataset] Ridley et al., 2019)
HadGEM3-GC31-HH 30km, 85L 7km, 75L ([dataset] Roberts, 2018)
CNRM-CM6-1 140km, 91L 70km, 75L ([dataset] Voldoire, 2019b)
CNRM-CM6-1-HR 50km, 91L 20km, 75L ([dataset] Voldoire, 2019a)
@@
Fig. S.1: Cross-shore transects of $VT_{\mbox{up-seas}}$ (color-filled black contours) for the Canary (top) and Benguela (bottom) upwelling system, within each sub-region (column). $VT_{\mbox{up-seas}}$ corresponds to an along-shore seasonal average over the entire historical period ($1995-2014$). The three-month seasonal average is adapted to each sub-region to only consider months most favourable to upwelling. Contours of potential temperature (light grey) and mixed layer depth (solid white) are superimposed. Line plots above each transect show cross-shore profiles of Sea Surface Temperature (solid black), offshore Ekman transport (solid red)and $VT_{m}$ (blue). The model shown here is the HadGEM3-GC3.1-HH model.
@@
Fig. S.2: Cross-shore transects of $VT_{\mbox{up-seas}}$ (color-filled black contours) for the Canary (top) and Benguela (bottom) upwelling system, within each sub-region (column). $VT_{\mbox{up-seas}}$ corresponds to an along-shore seasonal average over the entire historical period ($1995-2014$). The three-month seasonal average is adapted to each sub-region to only consider months most favourable to upwelling. Contours of potential temperature (light grey) and mixed layer depth (solid white) are superimposed. Line plots above each transect show cross-shore profiles of Sea Surface Temperature (solid black), offshore Ekman transport (solid red)and $VT_{m}$ (blue). The model shown here is the HadGEM3-GC3.1-LL model.
@@
Fig. S.3: Cross-shore transects of $VT_{\mbox{up-seas}}$ (color-filled black contours) for the Canary (top) and Benguela (bottom) upwelling system, within each sub-region (column). $VT_{\mbox{up-seas}}$ corresponds to an along-shore seasonal average over the entire historical period ($1995-2014$). The three-month seasonal average is adapted to each sub-region to only consider months most favourable to upwelling. Contours of potential temperature (light grey) and mixed layer depth (solid white) are superimposed. Line plots above each transect show cross-shore profiles of Sea Surface Temperature (solid black), offshore Ekman transport (solid red)and $VT_{m}$ (blue). The model shown here is the CNRM-CM6-1-HR model.
@@
Fig. S.4: Cross-shore transects of $VT_{\mbox{up-seas}}$ (color-filled black contours) for the Canary (top) and Benguela (bottom) upwelling system, within each sub-region (column). $VT_{\mbox{up-seas}}$ corresponds to an along-shore seasonal average over the entire historical period ($1995-2014$). The three-month seasonal average is adapted to each sub-region to only consider months most favourable to upwelling. Contours of potential temperature (light grey) and mixed layer depth (solid white) are superimposed. Line plots above each transect show cross-shore profiles of Sea Surface Temperature (solid black), offshore Ekman transport (solid red)and $VT_{m}$ (blue). The model shown here is the CNRM-CM6-1 model.
@@
Fig. S.5: Cross-shore transects of $VT_{\mbox{up-seas}}$ (color-filled black contours) for the Canary (top) and Benguela (bottom) upwelling system, within each sub-region (column). $VT_{\mbox{up-seas}}$ corresponds to an along-shore seasonal average over the entire future period ($2031-2050$). The three-month seasonal average is adapted to each sub-region to only consider months most favourable to upwelling. Contours of potential temperature (light grey) and mixed layer depth (solid white) are superimposed. Line plots above each transect show cross-shore profiles of Sea Surface Temperature (solid black), offshore Ekman transport (solid red)and $VT_{m}$ (blue). The model shown here is the HadGEM3-GC3.1-HH model.
@@
Fig. S.6: Cross-shore transects of $VT_{\mbox{up-seas}}$ (color-filled black contours) for the Canary (top) and Benguela (bottom) upwelling system, within each sub-region (column). $VT_{\mbox{up-seas}}$ corresponds to an along-shore seasonal average over the entire future period ($2081-2100$). The three-month seasonal average is adapted to each sub-region to only consider months most favourable to upwelling. Contours of potential temperature (light grey) and mixed layer depth (solid white) are superimposed. Line plots above each transect show cross-shore profiles of Sea Surface Temperature (solid black), offshore Ekman transport (solid red)and $VT_{m}$ (blue). The model shown here is the HadGEM3-GC3.1-MM model.
@@
Fig. S.7: Cross-shore transects of $VT_{\mbox{up-seas}}$ (color-filled black contours) for the Canary (top) and Benguela (bottom) upwelling system, within each sub-region (column). $VT_{\mbox{up-seas}}$ corresponds to an along-shore seasonal average over the entire future period ($2081-2100$). The three-month seasonal average is adapted to each sub-region to only consider months most favourable to upwelling. Contours of potential temperature (light grey) and mixed layer depth (solid white) are superimposed. Line plots above each transect show cross-shore profiles of Sea Surface Temperature (solid black), offshore Ekman transport (solid red)and $VT_{m}$ (blue). The model shown here is the HadGEM3-GC3.1-LL model.
@@
Fig. S.8: Cross-shore transects of $VT_{\mbox{up-seas}}$ (color-filled black contours) for the Canary (top) and Benguela (bottom) upwelling system, within each sub-region (column). $VT_{\mbox{up-seas}}$ corresponds to an along-shore seasonal average over the entire future period ($2081-2100$). The three-month seasonal average is adapted to each sub-region to only consider months most favourable to upwelling. Contours of potential temperature (light grey) and mixed layer depth (solid white) are superimposed. Line plots above each transect show cross-shore profiles of Sea Surface Temperature (solid black), offshore Ekman transport (solid red)and $VT_{m}$ (blue). The model shown here is the CNRM-CM6-1-HR model.